OpenStudy (anonymous):

f(x)=3xh(x) h(2)=-1 h ' (2)=-1 Find f ' (2)

OpenStudy (anonymous):

\[\bf f' = 3(xh)' = 3(x'h + xh')\]

OpenStudy (anonymous):

Soo f' = 3h + 3xh'

OpenStudy (anonymous):

so f'(2) =3h(2) + 3xh'(2) = ?....

OpenStudy (anonymous):

= -3 + 3*2*(-1) = -9

OpenStudy (anonymous):

Makes sense. Thanks, Mikael.

OpenStudy (anonymous):

tx